If X and Y are complementary angles, sinX =15/17 and cos X = 8/17, find each of the following :

a. tanX=

b. sinY=

c. cosY=

d. tsnY=


1
Expert's answer
2021-03-07T17:09:59-0500

"tan X = \frac{Sin X}{Cos X}"


"tan X = \frac{\frac{15}{17}}{\frac{8}{17}}"


"tan X = \frac{15}{17} × \frac{17}{8}"


"tan X = \frac{15}{8}"


Since X and Y are complementary angles


"Sin Y = Cos X = \frac{8}{17}"

"=>Sin Y = \frac{8}{17}"


"Cos Y = Sin X = \frac{15}{17}"

"=> Cos Y= \frac{15}{17}"


"tan Y = \frac{Sin Y}{Cos Y}"


"tan Y = \frac{\frac{8}{17}}{\frac{15}{17}}"


"=> tan Y = \frac{8}{17} × \frac{17}{15}"

"tan Y = \frac{8}{15}"